# A cutting-surface consensus approach for distributed robust optimization of multi-agent systems

### v2

@@ 6.2 Proof of Theorem 2
Suppose that the stepsize sequence $\left\{\alpha(t)\right\}$ is nonnegative and nonincreasing and Assumptions 2, 3, 4 hold. Define $\hat{\theta}_{i}(t)=\frac{\sum_{r=0}^{t}\alpha(r)\theta_{i}(r)}{\sum_{r=0}^{t}\alpha(r)}$, $\hat{w}(t)=\frac{\sum_{r=0}^{t}\alpha(r)w(r)}{\sum_{r=0}^{t}\alpha(r)}$. For the algorithm (3), the following holds:
@@
Proof 6.30.
We proceed in three steps: (i) Using the inequality (35) in Proposition 6.27 to give an upper estimate for the sum $\sum_{r=0}^{t}\left\{2\alpha(r)c^{\top}(w(t)-\theta^{*})+D_{4}\sum_{i=1}^{m}\|\phi_{i}(r)\|^{2}\right\}$ in terms of $\sum_{r=0}^{t}\alpha(r)$ and $\sum_{r=0}^{t}\alpha^{2}(r)$; (ii) Relating the left side of (43) to this sum by leveraging the convexity of $f_{0}(\theta)=c^{\top}\theta$ and Proposition 6.23; (iii) Proving (44) by using Lipschitz continuity of $f_{0}$ and (43).
@@
$\sum_{r=0}^{t}\bigg{[}\alpha(r)c^{\top}(w(r)-\theta^{*})+\frac{D_{4}}{2}\Phi(r)\bigg{]}\leq R_{1}+\frac{1}{2}\sum_{r=0}^{t}e(r)$ (46)
where $R_{1}=\frac{1}{2}\sum_{i=1}^{m}\pi_{i}(0)\|\theta_{i}(0)-\theta^{*}\|$. Next, we will establish an upper bound for the term $\sum_{r=0}^{t}e(r)$ from the following estimates:
$\sum_{r=0}^{t}\alpha(r)\lambda^{r}\leq\sum_{r=0}^{t}\alpha(0)\lambda^{r}\leq\frac{\alpha(0)}{1-\lambda},$ (47)
@@
$=\sum_{s=0}^{t-1}\alpha^{2}(s)\sum_{r=s+1}^{t}\lambda^{r-1-s}\leq\sum_{s=0}^{t-1}\frac{\alpha^{2}(s)}{1-\lambda},$
Hence, $\sum_{r=0}^{t}e(r)\leq\frac{D_{13}}{1-\lambda}\alpha(0)+(\frac{D_{23}}{1-\lambda}+D_{5})\sum_{r=0}^{t}\alpha^{2}(r)$. Therefore,
$\sum_{r=0}^{t}\!\bigg{[}\alpha(r)c^{\top}(w(r)\!-\!\theta^{*})\!+\!\frac{D_{4}}{2}\Phi(r)\bigg{]}\!\leq\!M_{1}\!+\!M_{2}\sum_{r=0}^{t}\alpha^{2}(r)$ (49)
@@
$\sum_{r=0}^{t}\psi(r)\leq\frac{1}{1-\lambda}\sum_{r=0}^{t-1}\bigg{(}\frac{\alpha^{2}(r)}{4}+\beta^{2}(r)\bigg{)}$ (57)
@@
where $C_{1}=M_{1}+M_{3}$, $C_{2}=M_{2}+M_{4}$. Dividing both sides by $\sum_{r=0}^{t}\alpha(r)$ and subsequently applying (50) and (51) leads to (43).
@@
$|c^{\top}(\hat{\theta}_{i}(t)-\theta^{*})|$ $\leq|c^{\top}(\hat{\theta}_{i}(t)-\hat{w}(t))|+c^{\top}(\hat{w}(t)-\theta^{*})$ (61)
$\leq C\|\hat{\theta}_{i}(t)-\hat{w}(t)\|+c^{\top}(\hat{w}(t)-\theta^{*})$
@@
According to Proposition 6.29, this result is straightforward. $\blacksquare$

### v3

@@ 5.2 Proof of Theorem 2
Suppose that the stepsize sequence $\left\{\alpha(t)\right\}$ is nonnegative and nonincreasing and Assumptions 24 hold. Define $\hat{\theta}_{i}(t)=\frac{\sum_{r=0}^{t}\alpha(r)\theta_{i}(r)}{\sum_{r=0}^{t}\alpha(r)}$, $\hat{w}(t)=\frac{\sum_{r=0}^{t}\alpha(r)w(r)}{\sum_{r=0}^{t}\alpha(r)}$. For the algorithm (3), the following holds:
@@
Proof 5.29.
We proceed in three steps: (i) Using Proposition 5.26 to give an upper estimate for the sum $\sum_{r=0}^{t}\{2\alpha(r)c^{\top}(w(r)-\theta^{*})+D_{4}\sum_{i=1}^{m}\|\phi_{i}(r)\|^{2}\}$ in terms of $\sum_{r=0}^{t}\alpha(r)$ and $\sum_{r=0}^{t}\alpha^{2}(r)$; (ii) Relating the left side of (41) to this sum by leveraging the convexity of $f_{0}$ and Proposition 5.22; (iii) Proving (42) by using Lipschitz continuity of $f_{0}$ and (41).
@@
$\sum_{r=0}^{t}\bigg[\alpha(r)c^{\top}(w(r)-\theta^{*})+\frac{D_{4}}{2}\Phi(r)\bigg]\leq R_{1}+\frac{1}{2}\sum_{r=0}^{t}e(r),$ (44)
where $R_{1}\!=\!\frac{1}{2}\sum_{i=1}^{m}\pi_{i}(0)\|\theta_{i}(0)\!-\!\theta^{*}\|$. Next, we will establish an upper bound for $\sum_{r=0}^{t}e(r)$ from the following estimates:
$\hskip-71.13188pt\sum_{r=0}^{t}\alpha(r)\lambda^{r}\leq\sum_{r=0}^{t}\alpha(0)\lambda^{r}\leq\frac{\alpha(0)}{1-\lambda},$ (45)
@@
$=\sum_{s=0}^{t-1}\alpha^{2}(s)\sum_{r=s+1}^{t}\lambda^{r-1-s}\leq\sum_{s=0}^{t-1}\frac{\alpha^{2}(s)}{1-\lambda}.$
Hence, $\sum_{r=0}^{t}e(r)\leq\frac{D_{13}}{1-\lambda}\alpha(0)+(\frac{D_{23}}{1-\lambda}+D_{5})\sum_{r=0}^{t}\alpha^{2}(r)$. Therefore, we have
$\hskip-5.69054pt\sum_{r=0}^{t}\!\bigg[\alpha(r)c^{\top}(w(r)-\theta^{*})\!+\!\frac{D_{4}}{2}\Phi(r)\bigg]\!\leq\!M_{1}\!+\!M_{2}\sum_{r=0}^{t}\alpha^{2}(r)$ (47)
@@
$\sum_{r=0}^{t}\psi(r)\leq\frac{1}{1-\lambda}\sum_{r=0}^{t-1}\bigg(\frac{\alpha^{2}(r)}{4}+\beta^{2}(r)\bigg).$ (55)
@@
where $C_{1}=M_{1}+M_{3}$, $C_{2}=M_{2}+M_{4}$. Dividing both sides by $\sum_{r=0}^{t}\alpha(r)$ and then applying (48) and (49) leads to (41).
@@
$\hskip-5.69054pt\begin{aligned} |c^{\top}(\hat{\theta}_{i}(t)-\theta^{*})|&\leq|c^{\top}(\hat{\theta}_{i}(t)-\hat{w}(t))|+c^{\top}(\hat{w}(t)-\theta^{*})\\ &\leq C\|\hat{\theta}_{i}(t)-\hat{w}(t)\|+c^{\top}(\hat{w}(t)-\theta^{*}).\end{aligned}$ (59)
@@
This result follows from Proposition 5.28. $\blacksquare$